ROTATIONAL SPECTROSCOPY AND MOLECULAR STRUCTURE OF ^{15}N_{2}-^{14}N_{2}O

HELEN O. LEUNG, Department of Chemistry, Mount Holyoke College, South Hadley, MA 01075-6407.

The rotational spectrum of ^{15}N_{2} −^{14} N_{2}O has been recorded in the 7-19 GHz region using a pulsed molecular beam, Fourier transform microwave spectrometer. Both a- and b-type transitions have been observed. The analysis of the hyperfine structure due to the two ^{14}N nuclei in the N_{2}O subunit reveals that the energy levels are doubled, owing to a tunneling motion of the ^{15}N_{2} subunit. The rotational constants support a planar, T-shaped structure, with ^{15}N_{2} forming the leg of the T. This geometry is consistent with that obtained using infrared spectroscopy. ^{a} The nuclear quadrupole coupling constants of the two ^{14}N nuclei indicate that the b-axis of the complex forms an angle between 10-12^{∘} with the N_{2}O axis.

aR. W. Randall, T. D. Dyke, and B. J. Howard, Faraday Discuss. Chem. Soc. \underline{86,} 21 (1988).